OpenStudy (anonymous):

The length of a rectangle is 2 less than3 times the width. The perimeter is 60 inches.Find the width.

OpenStudy (psymon):

We need to be able to translate the words given to us into mathematical operations and hopefully a full equation. So the question is giving us an equation for the length. L = something. So what is equals is 2 less than. The 2 less than means that L equals some number - 2. Furthermore, it says 2 less than 3 times the width. So some number - 2 = L, and that some number happens to be 3 times the width. We can write all of that out into an equation that would say 3w - 2 = L. Now we need a perimeter. The equation for the perimeter of a rectangle is 2L + 2w. We are being told that the perimeter is 60 inches, so essentially 2L + 2W = 60. So now we have two equations. These two equations can be solved in the same manner that you would for a system of equations. You can do substitution or elimination (addition) method. Since we are already told what L is, I will use substitution. Given that L = 3w-2, I will replace that into the 2nd equation. 2(3w - 2) + 2w = 60. So multiplying and then solving for w: 6w - 4 + 2w = 60. 8w - 4 = 60 and 8w = 64. Solving for w, w = 8. So the widh you need is 8.

OpenStudy (anonymous):

Thanks:)

OpenStudy (psymon):

Np ^_^